OpenStudy (anonymous):

Algebra2 The hyperbola defined by the equation below opens in which direction? Select all answers that apply. (y-6)^2 (x+5)^2 ________-________=1 8^2 2^2 Up Right Down LEFT

OpenStudy (anonymous):

Help

OpenStudy (anonymous):

anyyonee

OpenStudy (mertsj):

\[\frac{(x-h)^2}{a^2}-\frac{(y-k)^2}{b^2}=1\]

OpenStudy (mertsj):

That opens to the right. Now you can figure out which way yours opens.

OpenStudy (anonymous):

I dont understand

OpenStudy (anonymous):

anyoneee else care to explain or aanswer please??